Status Report

ISS MER Daily Report 18 Mar 2001

By SpaceRef Editor
March 18, 2001
Filed under ,









ISS MER Daily Report


Date: 3/18/01


STS 98/ISS 5A.1


GMT: 76:06:00 To GMT: 077:06:00


(12 a.m. 3-17-01 to 12:00 a.m. 3-18-01)


OVERALL SYSTEM STATUS:



ISS/Shuttle docked, ISS in control of attitude:




  • FGB: SD 1,3,4 and 5 powered off do to telemetry problems, all other SD on, Dew Point 52.8 deg. F, 11.5 deg C (52.8 deg. F, 11.5 deg C); EPS: Nominal; Battery 6 offline in deep discharge (nominal). FGB Komparus failed and deactivated. C&DH: FGB-2 MDM on, FGB-1 MDM off.

  • SM Elektron is off. Vozdukh is on. BMP status is on. CKB1 is on and in condensing mode; CKB2 is failed sensor and is non operational, powered off. Rapid depress response is off, Low-pressure warning is enabled. Dew Point 49.8 deg. F, 9.8 deg C (52.5 deg. F, 11.4 deg C). EPS: Nominal; Batteries in incomplete charge mode. SM ARCUs 22 online. SMTC: SMTC-1, 2 are operating in the redundant set. SMTC-3 was reported out of the set at GMT:2001/066:14:00 GMT; SMCC: SMCC 1, 2, and 3 are all operating in the redundant set. SMCC-1 is primary.

  • NODE1 ACS: Pressure limiting inhibited. Both Node 1 PPRV’s are not capped-MOD uncapped during PMA3 depress. IMV Fans: Node 1 Aft Port operating; all others off. IMV Valves: Node 1 aft port open and enabled; Aft Stbd closed and enabled; Port-fwd inhibited, Deck Aft is open and enabled; Deck Fwd is closed and enabled; Stbd both inhibited; Fwd both enabled and open. FDS: SD-1 enabled, SD-2 Smoke detector off due to RPC Trip. THC: Node 1 cabin fan is operating; RAMV (N1 side) is enabled in 90% cool for MPLM. C&DH: N1-2 MDM is primary, N1-1MDM is secondary

  • LAB PCA: Lab PCA on and monitoring. Positive Pressure relief is enabled VRCV and VRIV are closed. IMV Vlv/Fans: aft IMV valves are enabled and open. Aft port IMV fan is on. Fwd Stbd IMV valve is enabled and open. Fwd Stbd IMV fan is on. FDS: Smoke detectors LAB1PD1, LAB1SD5, & LAB1D6 on. CCAA: LAB1P6 full up. CDRA: Fail state. TCCSS: Full up state. MCA: Operational and providing US lab and Node. Vacuum System: VRS & VES are closed. All VES RIVs are closed. Water Venting: Post Heat State. Condensation tank quantity indicates 70.3 lbm (70.42 lbm). Condensate "T" is not installed. Cabin Pressure: 14.66 psia (14.65), Cabin Temp 74.3 deg. F. (76.1 deg. F), ppCO2 is 3.1 mmHg (3.9 mmHg), ppO2 is 162.2mmHg (158.7 mmHg). EPS: is nominal., RWS-Lab UOP 7 had GFCI trip during checkout. RPCM N13B-A RPCM 16 (Node SD2) experienced an over current trip; TCS: ATCS &PTCS are nominal;. ITCS: LT and MT in dual loop mode. The low temperature 3-way mixing valve is in full flow through, with outlet temperature of 44.8 deg. F (7.1 deg. C). The moderate temperature 3-way mixing valve is controlling to outlet temperature to 63.3 deg. F (17.4 deg. C.) Rack temperatures are nominal. C&T: Ecomm is powered off. S-band: configured for HDR; Audio Subsystem: configured for IAC2 and operations are nominal. AUAI2S, RAIU 1 have PBIT Faults. Video System: Operational. CBCS: Video cabin from ISS to PMA2 and shuttle in check out testing. C&DH: C&C MDMs: 1 is primary, 2 is backup, 3 is in standby. INT MDM: INT-2 is operating, INT-1 is off. GNC MDMS: GNC-1 primary, GNC-2 backup; LA MDMs: LA-1, LA-2, and LA-3 operating.; PL MDM: PL-1 primary, PL-2 Standby

  • Crew Systems: TVIS is operational with foot plate installed. IRED only one canister operational, no fix til 6A.

  • MPLM: Temperature is 75.38 deg. F, 24.1 deg C. (76.8 deg. F, 24.8 deg C). (MPLM shell heaters have been turned off). Overhead Aft IMV enabled and open; Overhead Fwd IMV valve closed; cabin fan on, smoke detector on; PPRA 1, 2, and 3 are closed. MDM Operating.

  • P6: Both PCUs in discharge mode. PVCU MDM 2B is prime/4B Backup. BGAs in Auto Track Mode. EETCS Loops A&B running on A pumps; FCV enabled; PFCS line heater 1&2 for loop B and B operational. RPCMs: N13B-A & N14B-C have health flags. PMCU: PMCU-1 MDM primary, MPCU-2 off.

  • GNC: ISS is currently flying in TEA ISS/CMG TA in control of Mated stack, Beta Angle —8.3 deg. Magnitude is decreasing (—15.9 degrees and mag decreasing). Avg. altitude is 213.3 x 204.9 nm 395.1 x 379.4 km (211.7 x 203 nm; 392.18 x 376.02 km). S&M subsystems quiescent and operable.









 


Previously Reported Status



  • Activation of RWS

  • MPLM transfer activities

  • Trouble shooting of CBCS

  • Node 1 8A & 9A heaters forced to enable. 2nd occurrence within 24 hours.


  • C&C #3 MDM successfully loaded and reinitialized.


 


Last 24 Hours Status



  • SD-2 RPCM trip during power up post charcoal filter replacement activities.

  • RWS Checkout tripped UOP 7. Reconfigured for one monitor only, checkout completed.

  • ISS reboost

  • MPM transfer activities.

  • ISS initialized vents (Pre heated to 260 deg. F) for impending water dump.


SIGNIFICANT EVENTS SCHEDULED FOR THE NEXT 24 HOURS:



  • Water dump 8 a.m.

  • MPLM Packing

  • MPLM Deactivation

  • Vestibule Depress

  • MPLM Re-Berthing operations.

 


 











































































































































Open Chits


Number (ISS)


Response Due (GMT)


Summary of Progress


(OPEN ISSMER CHITS — LATE ISSMER CHITS — NON ISSMER CHITS)


Chits without further MER Input required Chits that require MER Manager Action!


The following chits are all in Mission STS-098_5A


ISS0009


2001/037:00:00


MOD to IPRUSSIA, ISSMER — "5A Increment Full C&W Checkout Procedure"


3/9 — MER-M is not ready to close this CHIT — DO NOT CLOSE


ISS0126


2001/046:18:00


ISSMER to IPRUSSIA,MOD — "Russian Depress Sensor Status"


2/26 — MOD response is in, need response from Russians


ISS0199


2001/060:17:04


MOD to ISSMER — "S-bd Forward Link Audio Demux Activity"


3/9 — Expect info from HB and KSC today or tomorrow to close early next week. C&T Test scheduled for ISL on 3/22 to characterize data stream corruption of MRELP


ISS0204


2001/060:00:00


MOD to ISSMER — "Add NCS Loadshed Table 5 to Recon"


3/10 — ISSMER response provided.


ISS0206


2001/088:15:00


ISSMER to IPRUSSIA — "SMCC lane drop on GMT day 0505"

 

ISS0211


2001/061:20:00


Post-5A.1


MOD to ISSMER — "LA 1,3 LLA 11 Channel Anomalies"


3/10 — HSV says data will not be available until post-5A.1. C&DH has concurrence from ODIN that this is OK.


TI0107


???


ISSMER to ??? — Loss of S-Band Forward Audio Detection and Recovery"


3/10 — C&T still working this. When ready to open, transfer to a 5A.1 chit.


TI0249


2001/????


ISSMER to IMC, MOD, IPRUSSIA – "The Use of L200 Foam for Insulation Used around the SM Air Conditioner"


3/9 — Want to add to 5A.1 Increment Ops job jar. HOLD Opening until after 5A.1 Mission. Should also transfer to 5A.1 Database.

       

The following chits are all in Mission STS-102_5A.1


ISS0035


2001/071:20:30


IMC to IPRUSSIA, ISSMER, MOD — "REQUIREMENTS: Updates to Stage 5A.1 Priorities"


ISS MER Response Provided


ISS0043


2001/071/23:00


MOD to ISSMER — "Mass Properties Information for RSC-E Assembly & Config"


MER Response Complete


ISS0091


2001/075:17:34


IPRUSSIA to MOD — "S-Band timeline request"

 

ISS0100


2001/075:06:00


IMC to IPRUSSIA, ISSMER, MOD, CSR — "5A.1 ADD Docked Day to Flight for MPLM Stowage"


3/17 — MER Response Completed


ISS0105


2001/077:14:39


IPRUSSIA to MOD — "Onboard Control Events Protocol & Data on C&W signals appearance in USOS"

 

ISS0108


2001/076:02:00


IMC to IPRUSSIA, ISSMER, MER, MOD — "5A.1 Transfer: FD9, Russian Manifest/Transfer List"


3/17 — MER Response Provided


ISS0110


2001/077:18:07


ISSMER to MOD, IMC — "TVIS Imagery"

 

ISS0114


2001/100:16:00


MOD to ISSMER; Request patch for PMCU dump packet time


Assigned to C&DH


ISS0115


2001/090:16:00


MOD to ISSMER; Request fix for LA-1 MDM LLA 15 channel fails


Assigned to C&DH


ISS0116


2001/077:16:00


IMC to CSR, ISSMER, MER, MOD — "5A.1 Transfer: (Shuttle Middeck) FD9, US Manifest/Transfer List"


S&M , L&M has responded. Need to have schedule review.


ISS0118


2001/078:00:00


IMC to ISSMER, MOD — "USL Window Constraint"


Safety reviewing

       

TI0113


??????


IMC to MER; 5A.1 Transfer (MPLM): FD9, US Manifest/transfer list

 

TI0132

 

"Request to Inhibit HRDL Fault Safing"

 

TI0133


2001/079:18:00


ISSMER to MOD; Ku-Band Software Mask Verification

 


















































































Closed Chits


Number (ISS)


Response Due (GMT)


Summary of Progress


The following chits are all in Mission STS-098_5A


ISS0207


2001/068:15:00


ISSMER to MOD — "Node 1 MDM CPU utilization dump requests"


The following chits are all in Mission STS-102_5A1


ISS0011


2001/066:12:00


ISSMER to IPRUSSIA — "Request for SMCC Lane Change Information"


ISS0076


2001/073:16:00


MOD to ISSMER — "S-Band RFG Antenna Self Test Failure"


ISS0088


2001/074:16:00


IMC to CSR, ISSMER, MER, MOD — "5A.1 Transfer (MPLM): FD8, US Manifest/Transfer List"


ISS0090


2001/074:21:00


IMC to ISSMER, CSR, MOD — "Requirements: Request for MACE II data download to PCMCIA"


ISS0096


2001/075:18:00


ISSMER to MOD — "Label SWIS Network Control Unit (NCU)"


ISS0106


2001/077:00:00


ISSMER to MOD — "Request BDTs to verify health of SSU 2B"


ISS0112


2001/076:18:30


MER to MOD; Lab Condensate Dump Nozzle Preheat Time Out Workaround


ISS0104


2001/076:12:00


ISSMER to MOD — "Antenna Calibration to avoid SSU radiation"


ISS0103


2001/076:16:00


IMC to CSR, ISSMER, MER, MOD — "5A.1 Transfer (Shuttle Middeck): FD8, US Manifest/Transfer List"


ISS0111


2001/076:08:00


MOD to ISSMER — "MPLM Heater Activation Times"


ISS0101


2001/076:02:00


ISSMER to IMC, MOD, CSR, MER — "Return of Dessicant Bags"

     
     
     
     
     


 


















































































































































































Open Items for Investigation (IFIs)


Number


Title/Summary


Status


170


PCU Tubes never reached temp when heaters active


3/6 — Final solution in work. PRACA?


210


FPP Failure to Acquire Data


3/6 — Need data from Eliason


218


EVA Tool does not fit TCS Nonpropulsive Vent bolt heat


3/8 — Need to have details of EVA bolt resolution at KSC. Ready to close?


227


Burnt wires on veloergometer control box


3/6 — Need data filled in by EV&CS


229


HRDL MSD Inaccessible


3/6 — Need data filled in by C&DH


230


Node 1 PMA Duct Missing Parts


3/16 — Need to find out from crew how duct was reconfigured following the Soyuz relocation. Parts are on orbit. Has the crew found them yet?


235


Treadmill Slats Broken


3/16 — Need data filled in by EV&CS Have FIAR. Can close when IFI data filled in.


241


Station Vibration During TVIS Exercise


3/16 — Need data filled in by S&M or EV&CS? Waiting for IWIS data.


244


Periodic Annunciation of S-Band XPDR Non-critical Failure advisories


3/8 — Need Software PR. Needs to go to ASCB to get approval to write PR. C&T and C&DH working this.


245


S-Band BSP Loss of Forward link Audio Channel RELP Sync


3/7 — Close to S/W PR 2124? C&T confirming with MEIT data that this is the same problem.


247


Hitch Pin on WIF adapter Separated from Lanyard


3/16 — Have FIAR. EV&CS providing additional data for closure.


248


Audio Subsystem AUAI2S PBIT Indication


3/7 — Hardware is working. C&T needs additional data dumps to diagnose problem. Check PBIT chit.


249


C&C-2 MDM Switched to Diagnostic from Primary


3/8 — PPL uploaded. Still in work by ASCB?


253


SM KURS-P "Not Normal" Indication


SM Nadir Port Set 1 Self test failed twice. MCC-M plans to test on aft port to try isolate problem to the cable. SM Aft port test passed 3/16 — VI/Jack Bacon working this. Need to fill in IFI data.


254


BGA Observed vs Commanded Position (part 3)


3/16 — Non problem, but need NASA S/S Manager concurrence. Need Flt Rule CR to document reqt to lock BGAs?


255


OCA Intermittent Operation


3/7 – Need data and plan filled in by C&T


259


Echo on S-Band (Ground)


3/8 — Troubleshooting procedure uplinked to crew (OCA 1652). Troubleshooting planned post5A.1.


260


CCS-2 MSD Year incorrect for Filename : LDF.txt


3/7 – Need data filled in (from Funny Database?) by C&DH


261


S-BAND Gimbal Temperatures


3/16 — Have PR20449. Ready to close?


266


N1-2 MDM EEPROM Checksum Error


3/16 — MER C&DH working issue. Need SW PR.


270


4B BGA Latch #2 Failure to engage

 

273


4B BGA Motor Current Trip

 

276


Circles on Nadir CBM flange photographed

 

277


Center line Berthing system unusable

 

278


Internal Audio Controller (IAC) #1 Failover to IAC#2

 

281


DDCU LA3B Output Voltage Bias higher than expected with Converter off

 

283


S Band RFG Bit Summary

 

284


MPLM RPC-11 Output Voltage above expected value when off.

 

285


Node 1 8A Heater Forced to "Enable to Operate"

 

287


Rack K-Bar Fit Problems

 

288


UOP Trip During RWS Activation

 

289


N1-3B-A RPC 16 Overcurrent Trip

 

290


Lab Port water vent heaters did not achieve expected temperature.

 

291


S Band Radiation of STBD SSU

 

292


PCS Log Download Failure

 

294


Loss of PFR Attachment Device (PAD)


EV&CS will work this week of 3/19


295


PMA3 Connector P601 Unable to throw bail on stowage receptacle


EV&CS will enter data and complete week of 3/19


296


SM Gas Mask has 13 mm tear


ECLSS is working


297


Lab Fwd end cone Dzus fastener (Panel D0-D02)


PRACA generated; pending crew call down of panel removal and replace to ensure no further issue exists.


298


Low Water Vent Quantity

 

299


GLA’s turned off during Egress of MPLM

 

300


Damaged Threads on MPLM Passive alignment guide

 


 


































































Closed Items for Investigation (IFIs)


Number


Title/Summary


282


MTSAS Active Capture Latch Turn Count


264


RPC Tripped during LTA Activation


265


DZUS Fastener on Lab Panel 1S7_03


271


BGA Transition to Blind Mode Unexpectedly


275


P608 Dust Cap Cover


286


Node 1 3A Heater 3A & 8A Command Status Wrong


274


J603 O-Ring


198


SSC Network not Functioning


209


OCA Headset Cable Failure


251


SKV 2 Shutdown Due to Oil Pressure Sensor


252


EEATCS Loop A Radiator Return Temperature Spikes


257


Unexpected momentum/attitude disturbances


262


IRED Scraping


263


Seat Track Fit Problem


279


MPLM ELPS Current


280


N1N/MPLM Vestibule Closeout


293


Lab Condensate Tank Leak fail indication


 






















Action Items


Status

SpaceRef staff editor.